OpenStudy (tacotime):

Two point charges are placed on the x-axis as follows: charge q1 = 3.98 nC is located at x= 0.199 m , and charge q2 = 5.00 nC is at x= -0.295 m . What is the magnitude of the total force exerted by these two charges on a negative point charge q3 = -5.97 nC that is placed at the origin?

OpenStudy (tacotime):

I tried to do F1 = k*q1*q3/r^2 and F1 = k*q2*q3/r^2 then F1 + F2 but for some reason its wrong.

OpenStudy (michele_laino):

the requested force (magnitude with sign), is given by the subsequent formula: \[\Large F = - \frac{1}{{4\pi {\varepsilon _0}}}\frac{{\left| {{q_2}{q_3}} \right|}}{{d_2^2}} + \frac{1}{{4\pi {\varepsilon _0}}}\frac{{\left| {{q_1}{q_3}} \right|}}{{d_1^2}}\]